# Shuangrui Ding1  Rui Qian1  Xiaoyi Dong1,2  Pan Zhang2 Yuhang Zang2  Yuhang Cao2  Yuwei Guo1   Dahua Lin1   Jiaqi Wang2

### v2

@@ Abstract
The Segment Anything Model 2 (SAM 2) has emerged as a powerful foundation model for object segmentation in both images and videos, paving the way for various downstream video applications. The crucial design of SAM 2 for video segmentation is its memory module, which prompts object-aware memories from previous frames for current frame prediction. However, its greedy-selection memory design suffers from the “error accumulation” problem, where an errored or missed mask will cascade and influence the segmentation of the subsequent frames, which limits the performance of SAM 2 toward complex long-term videos. To this end, we introduce SAM2Long, an improved training-free video object segmentation strategy, which considers the segmentation uncertainty within each frame and chooses the video-level optimal results from multiple segmentation pathways in a constrained tree search manner. In practice, we maintain a fixed number of segmentation pathways throughout the video. For each frame, multiple masks are proposed based on the existing pathways, creating various candidate branches. We then select the same fixed number of branches with higher cumulative scores as the new pathways for the next frame. After processing the final frame, the pathway with the highest cumulative score is chosen as the final segmentation result. Benefiting from its heuristic search design, SAM2Long is robust toward occlusions and object reappearances, and can effectively segment and track objects for complex long-term videos. Without introducing any additional parameters or further training, SAM2Long significantly and consistently outperforms SAM 2 on six VOS benchmarks. Notably, SAM2Long achieves an average improvement of 3.0 points across all 24 head-to-head comparisons, with gains of up to 5.3 points in $\mathcal{J}\&\mathcal{F}$ on long-term video object segmentation benchmarks such as SA-V and LVOS. The code is released at https://github.com/Mark12Ding/SAM2Long.
@@ 1 Introduction
The Segment Anything Model 2 (SAM 2) has gained significant attention as a unified foundational model for promptable object segmentation in both images and videos. Notably, SAM 2 (Ravi et al., 2024) has achieved state-of-the-art performance across various video object segmentation tasks, significantly surpassing previous methods. Building upon the original SAM (Kirillov et al., 2023), SAM 2 incorporates a memory module that enables it to generate masklet predictions using stored memory contexts from previously observed frames. This module allows SAM 2 to seamlessly extend SAM into the video domain, processing video frames sequentially, attending to the prior memories of the target object, and maintaining object coherence over time.
@@
Figure 1: Comparison of occlusion handling and long-term capability between SAM 2 and SAM2Long. (a) When an occlusion occurs, SAM 2 may lose track or follow the wrong object, leading to accumulated errors. In contrast, SAM2Long utilizes memory tree search to recover when the object reappears. (b) The per-frame $\mathcal{J}\&\mathcal{F}$ scores of the predicted masks are plotted at specific timestamps on the LVOS and SA-V datasets. SAM2Long demonstrates greater resilience to elapsed time compared to SAM 2, maintaining superior performance over longer periods.
@@
While SAM 2 demonstrates strong performance in video segmentation, its greedy segmentation strategy struggles to handle complex video scenarios with frequent occlusions and object reappearance. In detail, SAM 2 confidently and accurately segments frames when clear visual cues are present. However, in scenarios with occlusions or reappearing objects, it can produce mask proposals that are highly variable and uncertain. Regardless of the frame’s complexity, a uniform greedy selection strategy is applied to both scenarios: the mask with the highest predicted IoU is selected. Such greedy choice works well for the easy cases but raises the error potential for the challenging frames. Once an incorrect mask is selected into memory, it is uncorrectable and will mislead the segmentation of the subsequent frames. We show such an “error accumulation” problem in Figure 1 both qualitatively and quantitatively. The performance of SAM 2 progressively deteriorates as the propagation extends into the later temporal segment, highlighting its limitations in maintaining accurate tracking over time.
To this end, we redesign the memory module of SAM 2 to enhance its long-term capability and robustness against occlusions and error propagation. Our improvement is completely free of additional training and does not introduce any external parameters, but simply unleashes the potential of SAM 2 itself. Our approach is motivated by the observation that the SAM 2 mask decoder generates multiple diverse masks, accompanied by predicted IoU scores and an occlusion score when handling challenging and ambiguous cases. However, SAM 2 only selects a single mask as memory, sometimes disregarding the correct one. To address this, we aim to equip SAM 2 with multiple memory pathways, allowing various masks to be stored as memory at each time step, thereby improving predictions for subsequent frames.
@@
In particular, we introduce a novel constrained tree memory structure, which maintains a fixed number of memory pathways over time to explore multiple segmentation hypotheses with efficiently managed computational resources. At each time step, based on a set of memory pathways, each with its own memory bank and cumulative score (accumulated logarithm of the predicted IoU scores across the pathway), we produce multiple candidate branches for the current frame. Then, among all the branches, we select the same fixed number of branches with higher cumulative scores and prune other branches, thereby constraining the tree’s growth. After processing the final frame, the pathway with the highest cumulative score is selected as the final segmentation result. Moreover, to prevent premature convergence on incorrect predictions, we select hypotheses with distinct predicted masks when their occlusion scores indicate uncertainty, in order to maintain diversity in the tree branches. This tree-like memory structure augments SAM 2’s ability to effectively overcome error accumulation.
@@
Within each pathway, we construct an object-aware memory bank that selectively includes frames with confidently detected objects and high-quality segmentation masks, based on the predicted occlusion scores and IoU scores. Instead of simply storing the nearest frames as SAM 2 does, we filter out frames where the object may be occluded or poorly segmented. This ensures that the memory bank provides effective object cues for the current frame’s segmentation. Additionally, we modulate the memory attention calculation by weighting memory entries according to their occlusion scores, emphasizing more reliable entries during cross-attention. These strategies help SAM 2 focus on reliable object clues and improve segmentation accuracy with negligible computational overhead. As evidenced in Figure 1(a), our approach successfully resolves occlusions and re-tracks the recurring balloon, where SAM 2 fails.
@@
We provide a comprehensive evaluation demonstrating that SAM2Long consistently outperforms SAM 2 across six VOS benchmarks, particularly excelling in long-term and occlusion-heavy scenarios. For instance, on the challenging SA-V test set, SAM2Long-L improves the $\mathcal{J}\&\mathcal{F}$ score by 5.3 points, and SAM2Long-S shows an impressive 4.7-point gain over the same size SAM 2 model on SA-V val set. Similar trends are observed on the LVOS validation set, where SAM2Long-S surpasses SAM 2-S by 3.5 points. These consistent improvements across different model sizes, including both SAM 2 and the more recent SAM 2.1 model weights, clearly demonstrate the effectiveness of our proposed method. Furthermore, as illustrated in Figure 1(b), the per-frame performance gap between SAM2Long and SAM 2 widens over time, showcasing SAM2Long excels in long-term tracking scenarios. With these results, we believe SAM2Long sets a new standard for video object segmentation based on SAM 2 in complex, real-world applications, delivering superior performance without any additional training or external parameters.
@@ 2.1 Video Object Segmentation
Perceiving the environment in terms of objects is a fundamental cognitive ability of humans. In computer vision, Video Object Segmentation (VOS) tasks aim to replicate this capability by requiring models to segment and track specified objects within video sequences. A substantial amount of research has been conducted on video object segmentation in recent decades (Fan et al., 2019a; Oh et al., 2019; Hu et al., 2018a; Oh et al., 2018; Perazzi et al., 2017; Wang et al., 2019; Hu et al., 2018b; Li & Loy, 2018; Bao et al., 2018; Zhang et al., 2019; Li et al., 2020; Johnander et al., 2019; Zhang et al., 2023; Ventura et al., 2019; Li et al., 2022; Wu et al., 2023b; Wang et al., 2023; Qian et al., 2023; 2025; Ding et al., 2022; 2023b).
There are two main protocols for evaluating VOS models (Pont-Tuset et al., 2017; Perazzi et al., 2016): semi-supervised and unsupervised video object segmentation. In semi-supervised VOS, the first-frame mask of the objects of interest is provided, and the model tracks these objects in subsequent frames. In unsupervised VOS, the model directly segments the most salient objects from the background without any reference. It is important to note that these protocols are defined in the inference phase, and VOS methods can leverage ground truth annotations during the training stage.
In this paper, we explore SAM 2 (Ravi et al., 2024), for its application in semi-supervised VOS. We enhance the memory design of SAM 2, significantly improving mask propagation performance without requiring any additional training.
@@ 2.2 Memory-Based VOS
Video object segmentation remains an unsolved challenge due to the inherent complexity of video scenes. Objects in videos can undergo deformation (Tokmakov et al., 2023), exhibit dynamic motion (Brox & Malik, 2010), reappear over long durations (Hong et al., 2024; 2023), and experience occlusion (Ding et al., 2023a), among other challenges. To address the above challenges, adopting a memory architecture to store the object information from past frames is indispensable for accurately tracking objects in video. Previous methods (Bhat et al., 2020; Caelles et al., 2017; Maninis et al., 2018; Robinson et al., 2020; Voigtlaender & Leibe, 2017) treat VOS as an online learning task, where networks are test-time tuned on the first-frame annotation. However, this approach was time-consuming due to test-time fine-tuning. Other techniques (Chen et al., 2018; Hu et al., 2018b; Voigtlaender et al., 2019; Yang et al., 2018; 2020; 2021b) use template matching, but they lack the capability of tracking under occlusion.
More recent approaches have introduced efficient memory reading mechanisms, utilizing either pixel-level attention (Cheng et al., 2023; Zhou et al., 2024; Duke et al., 2021; Liang et al., 2020; Oh et al., 2018; Seong et al., 2020; Cheng & Schwing, 2022; Xie et al., 2021; Yang & Yang, 2022; Yang et al., 2021a) or object-level attention (Athar et al., 2023; 2022; Cheng et al., 2024). A prominent example is XMem (Cheng & Schwing, 2022), which leverages a hierarchical memory structure for pixel-level memory reading combined. Building on XMem’s framework, Cutie (Cheng et al., 2024) further improves segmentation accuracy by processing pixel features at the object level to better handle complex scenarios.
The latest SAM 2 (Ravi et al., 2024) incorporates a simple memory module on top of the image-based SAM (Kirillov et al., 2023), enabling it to function for VOS tasks. However, by selecting only the temporally nearest frames as memory, SAM 2 struggles with challenging cases involving long-term reappearing objects and confusingly similar objects. we redesign SAM 2’s memory to maintain multiple potential correct masks, making the model more object-aware and robust.
2.3 Segment Anything Model
Segment Anything Model (SAM) (Kirillov et al., 2023) is recognized as a milestone vision foundation model that can segment any object in an image using interactive prompts. Its impressive zero-shot transfer performance has shown great versatility in various vision tasks, including segmentation applications (Li et al., 2023; Ma et al., 2024; Xu et al., 2024), image editing (Gao et al., 2023a) and object reconstruction (Lin et al., 2024).
Building on SAM, SAM 2 (Ravi et al., 2024) extends its functionality to video segmentation through a memory-based transformer architecture for real-time video processing. SAM 2’s memory stores information about objects and past interactions, enabling it to generate segmentation masks across video frames more accurately and efficiently than previous methods. To further enhance SAM 2, we introduce a constrained memory tree structure. This training-free design leverages the SAM 2’s ability to generate multiple candidate mask proposals with predicted IoU and occlusion score, mitigating error accumulation during segmentation.
@@ 3.1 Preliminary on SAM 2
SAM 2 (Ravi et al., 2024) begins with an image encoder that encodes each input frame into embeddings. In contrast to SAM, where frame embeddings are fed directly into the mask decoder, SAM 2 incorporates a memory module that conditions the current frame’s features on both previous and prompted frames. Specifically, for semi-supervised video object segmentation tasks, SAM 2 maintains a memory bank at each time step $t\geq 1$:
@@ 3.2 Constrained Tree Memory with Uncertainty Handling
To enhance SAM 2’s robustness towards long-term and ambiguous cases, we propose a constrained tree memory structure that enables the model to explore various object states over time with minimal computational overhead. We show the high-level pipeline in Figure 2. This tree-based approach maintains multiple plausible pathways and mitigates the effects of occlusions and erroneous predictions.
@@
Compared to SAM 2, our approach introduces additional computation mainly by increasing the number of passes through the mask decoder and memory module. Notably, these components are lightweight relative to the image encoder. For instance, the image encoder of SAM 2-Large consists of $212$M parameters while the total parameter of SAM 2-Large is $224$M. Since we process the image encoder only once just as SAM 2 does, the introduction of a memory tree adds negligible computational cost while significantly enhancing SAM 2’s robustness against error-prone cases.
Uncertainty Handling. Unfortunately, there are times when all pathways are uncertain. To prevent the model from improperly converging on incorrect predictions, we implement a strategy to maintain diversity among the pathways by deliberately selecting distinct masks. That is, if the maximum absolute occlusion score across all pathways at time $t$, $\max(\{|o_{t}^{p}|\}_{p=1}^{P})$, is less than a predefined uncertainty threshold $\delta_{\text{conf}}$, we enforce the model to select mask candidates with unique IoU values. This is inspired by the observation that, within the same frame, different IoU scores typically correspond to distinct masks. In practice, we round each IoU score $\text{IoU}_{t}^{p,k}$ to two decimal places and only select those hypotheses with distinct rounded values.

### v3

# SAM2Long: Enhancing SAM 2 for Long Video Segmentation with a Training-Free Memory Tree

@@ Abstract
The Segment Anything Model 2 (SAM 2) has emerged as a powerful foundation model for object segmentation in both images and videos. The crucial design of SAM 2 for video segmentation is its memory module, which prompts object-aware memories from previous frames for current frame prediction. However, its greedy-selection memory design suffers from the “error accumulation” problem, where an errored or missed mask will cascade and influence the segmentation of the subsequent frames, which limits the performance of SAM 2 toward complex long-term videos. To this end, we introduce SAM2Long, an improved training-free video object segmentation strategy, which considers the segmentation uncertainty within each frame and chooses the video-level optimal results from multiple segmentation pathways in a constrained tree search manner. In practice, we maintain a fixed number of segmentation pathways throughout the video. For each frame, multiple masks are proposed based on the existing pathways, creating various candidate branches. We then select the same fixed number of branches with higher cumulative scores as the new pathways for the next frame. After processing the final frame, the pathway with the highest cumulative score is chosen as the final segmentation result. Benefiting from its heuristic search design, SAM2Long is robust toward occlusions and object reappearances, and can effectively segment and track objects for complex long-term videos. Without further training, SAM2Long significantly and consistently outperforms SAM 2 on nine VOS benchmarks and three VOT benchmarks. Notably, SAM2Long achieves an average improvement of 3.7 points across all 12 direct comparisons, with gains of up to 5.3 points in $&\mathcal{J}\&\mathcal{F}$ on long-term video object segmentation benchmarks such as SA-V and LVOS. The code is released at https://github.com/Mark12Ding/SAM2Long.
@@ 1 Introduction
The Segment Anything Model 2 (SAM 2) has gained significant attention as a unified foundational model for promptable object segmentation in both images and videos. Notably, SAM 2 [50] has achieved state-of-the-art performance across various video object segmentation tasks, significantly surpassing previous methods. Building upon the original SAM [33], SAM 2 incorporates a memory module that enables it to generate masklet predictions using stored memory contexts from previously observed frames. This module allows SAM 2 to seamlessly extend SAM into the video domain, processing video frames sequentially, attending to the prior memories of the target object, and maintaining object coherence over time.
@@
Figure 1: Comparison of occlusion handling and long-term capability between SAM 2 and SAM2Long. (a) When an occlusion occurs, SAM 2 may lose track or follow the wrong object, leading to accumulated errors. In contrast, SAM2Long utilizes memory tree search to recover when the object reappears. (b) Per-frame $&\mathcal{J}\&\mathcal{F}$ scores of the predicted masks with different timestamp are plotted on the LVOS and SA-V datasets. SAM2Long demonstrates greater resilience to long video compared to SAM 2.
@@
While SAM 2 demonstrates strong performance in video segmentation and downstream tasks [25], its greedy segmentation strategy struggles to handle complex video scenarios with frequent occlusions and object reappearance. In detail, SAM 2 confidently and accurately segments frames when clear visual cues are present. However, in scenarios with occlusions or reappearing objects, it can produce mask proposals that are highly variable and uncertain. Regardless of the frame’s complexity, a uniform greedy selection strategy is applied to both scenarios: the mask with the highest predicted IoU is selected. Such greedy choice works well for the easy cases but raises the error potential for the challenging frames. Once an incorrect mask is selected into memory, it is uncorrectable and will mislead the segmentation of the subsequent frames. Figure 1 illustrates the problem of “error accumulation” both qualitatively and quantitatively. The performance of SAM 2 gradually declines as the propagation progresses into later temporal segments, underscoring its limitations in keeping accurate tracking over time.
@@
To this end, we redesign the memory module of SAM 2 to enhance its long-term capability and robustness against occlusions and error propagation. Our approach is motivated by the observation that the SAM 2 mask decoder generates multiple diverse masks, accompanied by predicted IoU scores and an occlusion score when handling challenging and ambiguous cases. However, SAM 2 only selects a single mask as memory, sometimes disregarding the correct one. To address this limitation, we propose integrating SAM 2 with multiple memory pathways, inspired by prior Multiple Hypothesis Tracking (MHT) techniques [15, 32, 51]. These techniques, widely adopted in the tracking community, enable the storage of various masks as memory at each time step and delay the association decision to obtain video-level optimal result. In particular, we maintain a fixed number of memory pathways over time to explore multiple segmentation hypotheses with efficiently managed computational resources. At each time step, based on a set of memory pathways, each with its own memory bank and cumulative score (accumulated logarithm of the predicted IoU scores across the pathway), we produce multiple candidate branches for the current frame. Then, among all the branches, we only keep the same number of branches with higher cumulative scores and prune other branches, thereby constraining the tree’s growth. After processing the final frame, the pathway with the highest cumulative score is selected as the final segmentation result. Moreover, to prevent premature convergence on incorrect predictions, we select pathways with distinct predicted masks when their occlusion scores indicate uncertainty. In this way, we maintain diversity in the tree branches in the challenging and distraction scenario. This tree-like memory structure augments SAM 2’s ability to effectively overcome error accumulation.
@@
Within each pathway, we construct an object-aware memory bank that selectively includes frames with confidently detected objects and high-quality segmentation masks, based on the predicted occlusion scores and IoU scores. Instead of simply storing the nearest frames as SAM 2 does, we filter out frames where the object may be occluded or poorly segmented. This ensures that the memory bank provides effective object cues for the current frame’s segmentation. Additionally, we modulate the memory attention calculation by weighting memory entries according to their occlusion scores, emphasizing more reliable entries during cross-attention. These strategies help SAM 2 focus on reliable object clues and improve segmentation accuracy with negligible computational overhead. As evidenced in Figure 1(a), our approach successfully resolves occlusions and re-tracks the recurring dancers, where SAM 2 fails.
@@
Our improvement is completely free of additional training and does not introduce any external parameters, but simply unleashes the potential of SAM 2 itself. We provide a comprehensive evaluation that SAM2Long consistently outperforms SAM 2 across nine VOS benchmarks and three VOT benchmarks, particularly excelling in long-term and occlusion-heavy scenarios. Moreover, datasets with longer video durations generally lead to greater performance gains, which aligns with the motivation of our design. For instance, on the challenging SA-V test set, SAM2Long-L improves the $&\mathcal{J}\&\mathcal{F}$ score by 5.3 points, and SAM2Long-S shows an impressive 4.7-point gain over the same size SAM 2 model on SA-V val set. Similar trends are observed on the LVOS validation set, where SAM2Long-S surpasses SAM 2-S by 3.5 points. These consistent improvements across different model sizes, including both SAM 2 and the more recent SAM 2.1 model weights, clearly demonstrate the generalization ability of our proposed method. Furthermore, as shown in Figures 1(b), the per-frame performance gap between SAM2Long and SAM 2 increases over time. This highlights SAM2Long’s strength in long-term tracking scenarios. With these results, we believe SAM2Long sets a new standard for video object segmentation based on SAM 2 in complex, real-world applications.
@@ 2.1 Video Object Segmentation
Perceiving the environment in terms of objects is a fundamental cognitive ability of humans. In computer vision, Video Object Segmentation (VOS) tasks aim to replicate this capability by requiring models to segment and track specified objects within video sequences. A substantial amount of research has been conducted on video object segmentation in recent decades [21, 44, 28, 43, 46, 59, 29, 35, 3, 71, 36, 31, 70, 55, 34, 61, 58, 48, 49, 18, 19]. There are two main protocols for evaluating VOS models [47, 45]: semi-supervised and unsupervised video object segmentation. In semi-supervised VOS, the first-frame mask of the objects of interest is provided, and the model tracks these objects in subsequent frames. In unsupervised VOS, the model directly segments the most salient objects from the background without any reference. It is important to note that these protocols are defined in the inference phase, and VOS methods can leverage ground truth annotations during the training stage. In this paper, we explore SAM 2 [50], for its application in semi-supervised VOS. We enhance the memory design of SAM 2, significantly improving mask propagation performance without any additional training.
@@ 2.2 Memory-Based VOS
Video object segmentation remains an unsolved challenge due to the inherent complexity of video scenes. Objects in videos can undergo deformation [54], exhibit dynamic motion [6], reappear over long durations [27, 26], and experience occlusion [17], among other challenges. To address the above challenges, adopting a memory architecture to store the object information from past frames is indispensable for accurately tracking objects in video [5, 7, 40, 52, 56, 10, 29, 57, 64, 66, 68]. Recent approaches have introduced efficient memory reading mechanisms, utilizing either pixel-level attention [13, 73, 20, 37, 43, 53, 11, 62, 65, 67] or object-level attention [2, 1, 14]. Among these, Multiple Hypothesis Tracking (MHT) [15, 32, 51] stands out as a robust paradigm that maintains tree search to construct and manage memory. A prominent example is XMem [11], which leverages a hierarchical memory structure for pixel-level memory reading combined. Building on XMem’s framework, Cutie [14] further improves segmentation accuracy by processing pixel features at the object level to better handle complex scenarios. The latest SAM 2 [50] incorporates a simple memory module on top of the image-based SAM [33], enabling it to function for VOS tasks. However, SAM 2 struggles with challenging cases involving long-term reappearing objects and confusingly similar objects. Motivated by MHT, we redesign SAM 2’s memory to maintain multiple potential correct masks, making the model more object-aware and robust.
@@ 3.1 Preliminary on SAM 2
SAM 2 [50] begins with an image encoder that encodes each input frame into embeddings. In contrast to SAM, where frame embeddings are fed directly into the mask decoder, SAM 2 incorporates a memory module that conditions the current frame’s features on both previous and prompted frames. Specifically, for semi-supervised video object segmentation tasks, SAM 2 maintains a memory bank at each time step $t\geq 1$:
@@ 3.2 Constrained Tree Memory with Uncertainty Handling
To enhance SAM 2’s robustness towards long-term and ambiguous cases, we propose a constrained tree memory structure that enables the model to explore various object states over time with minimal computational overhead. We show the high-level pipeline in Figure 2. Note that we store memory separately for each target. Each target maintains its own individual tree memory bank. This tree-based approach maintains multiple plausible pathways and mitigates the effects of erroneous predictions.
@@
Compared to SAM 2, our approach introduces additional computation mainly by increasing the number of passes through the mask decoder and memory module. Notably, these components are lightweight relative to the image encoder. For instance, the image encoder of SAM 2-Large consists of 212212212M parameters while the total parameter of SAM 2-Large is 224224224M. Since we process the image encoder only once just as SAM 2 does, the introduction of a memory tree adds negligible computational cost while significantly enhancing SAM 2’s robustness against error-prone cases.
Method SA-V val SA-V test LVOS v2 val
[…]
SAM2-T 73.5 70.1 76.9 74.6 71.1 78.0 77.8 74.5 81.2
SAM2.1-T 75.1 71.6 78.6 76.3 72.7 79.8 81.6 77.9 85.2
SAM2-S 73.0 69.7 76.3 74.6 71.0 78.1 79.7 76.2 83.3
SAM2.1-S 76.9 73.5 80.3 76.9 73.3 80.5 82.1 78.6 85.6
SAM2-B+ 75.4 71.9 78.8 74.6 71.2 78.1 80.2 76.8 83.6
SAM2.1-B+ 78.0 74.6 81.5 77.7 74.2 81.2 83.1 79.6 86.5
SAM2-L 76.3 73.0 79.5 75.5 72.2 78.9 83.0 79.6 86.4
SAM2Long-L 80.8 (4.5$\uparrow$) 77.1 84.5 80.8 (5.3$\uparrow$) 76.8 84.7 85.4 (2.4$\uparrow$) 81.8 88.7
SAM2.1-L 78.6 75.1 82.0 79.6 76.1 83.2 84.0 80.7 87.4
Table 1: Performance comparison on SA-V [50] and LVOS v2 [27] datasets between SAM 2 and SAM2Long across all model sizes. $\dagger$ We report the re-produced performance of SAM 2(2.1) using its open-source code and checkpoint.
Uncertainty Handling. Unfortunately, there are times when all pathways are uncertain. To prevent the model from improperly converging on incorrect predictions, we implement a strategy to maintain diversity among the pathways by deliberately selecting distinct masks. That is, if the maximum absolute occlusion score across all pathways at time $t$, $\max(\{|o_{t}^{p}|\}_{p=1}^{P})$, is less than a predefined uncertainty threshold $\delta_{\text{conf}}$, we enforce the model to select mask candidates with unique IoU values. This design is inspired by the observation that, within the same frame, different IoU scores typically correspond to distinct masks, which we verify in Table 8. In practice, we round each IoU score $\text{IoU}_{t}^{p,k}$ to two decimal places and only select those hypotheses with distinct rounded values.